# Comparative Study of State-based Neural Networks for Virtual Analog Audio Effects Modeling

### v5

@@ Abstract
Analog electronic circuits are at the core of an important category of musical devices, which includes a broad range of sound synthesizers and audio effects. The nonlinear features of their passive and active electronic components give analog musical devices a distinctive timbre and sound quality, making them highly desirable. The development of software that simulates analog musical devices, known as virtual analog modeling, is a significant sub-field in audio signal processing. Artificial neural networks are a promising technique for virtual analog modeling. They have rapidly gained popularity for the emulation of analog audio effects, particularly recurrent networks. While neural approaches have been successful in accurately modeling distortion circuits, they require architectural improvements that account for parameter conditioning and low-latency response. Although hybrid solutions can offer advantages, black-box approaches can still be advantageous in some contexts. In this article, we explore the application of recent machine learning advancements for virtual analog modeling. In particular, we compare State-Space models and Linear Recurrent Units against the more common Long Short-Term Memory networks. These have shown promising ability in sequence-to-sequence modeling tasks, showing a notable improvement in signal history encoding. Our comparative study uses these black-box neural modeling techniques with a variety of audio effects. We evaluate the performance and limitations of these models using multiple metrics, providing insights for future research and development. Our metrics aim to assess the models’ ability to accurately replicate energy envelopes and frequency contents, with a particular focus on transients in the audio signal. To incorporate control parameters into the models, we employ the Feature-wise Linear Modulation method. Long Short-Term Memory networks exhibit better accuracy in emulating distortions and equalizers, while the State-Space model, followed by Long Short-Term Memory networks when integrated in an encoder-decoder structure, and Linear Recurrent Unit outperforms others in emulating saturation and compression. When considering long time-variant characteristics, the State-Space model demonstrates the greatest capability to track history. The Long Short-Term Memory networks tend to introduce audio artifacts.
@@ 1 Introduction
The emulation of analog musical devices has become pervasive in digital products due to their distinct sound, appealing to producers and musicians. Analog circuits use components like operational amplifiers and diodes, whose nonlinear behaviors contribute to creating a unique sound. Replicating these devices, called Virtual Analog (VA) modeling, is now a crucial area in digital audio signal processing.
@@
There are two possible approaches for emulating analog audio effects: physics-based and data-driven. Physics-based approaches discretize existing continuous-time models. Data-driven approaches identify a discrete model directly from discrete measurement data. Machine learning, as a data-driven method, has become widely used in audio modeling and is often integrated into more traditional digital signal-processing frameworks. Hybrid solutions, incorporating knowledge of the modeling tasks, also exist and can offer advantages, but data-driven approaches can still be appealing in some contexts. While the former may not always be accurate enough, the latter may not be feasible for certain tasks.
@@
Data-driven models, also called Black-box, being applicable to different audio modeling tasks, have already been proposed  (Ramírez et al.,, 2020a), but these usually involve large networks that are not suitable for low-latency and real-time applications. In addition, while machine learning modeling of nearly all types of analog effects has been investigated, they often do not provide a satisfactory response when continuously manipulating conditioning parameters, which is needed to offer users runtime control over the effect’s parameters.
This article presents a comparative study that applies recurrent and state-based neural network architectures to model audio effects with different complexities. Neural networks utilizing internal states and a recurrent formulation are often preferred among the different types of artificial neural networks. The most known type is the Recurrent Neural Networks (RNNs). A Similar formulation is present in the State-Space Networks (SSMs). The internal states function as memory, making them particularly suitable in scenarios where time dependencies are crucial. Especially when the inputs are limited to one or a small segment of digital audio samples. For this reason, in this study, we consider Convolutional Neural Networks (CNNs) (LeCun et al.,, 1998) only when integrated with RNNs  (Simionato & Fasciani,, 2022). The aim is to work with small input vectors while also considering long-term temporal dependencies. Therefore, we focus on state-based structures. Analog audio effects are dynamic physical systems that exhibit time-dependent characteristics: their output depends on previous internal states. Therefore, dynamic physical systems possess internal states that describe their physical evolution, similar to stateful machine learning algorithms.
@@
We designed five architectures based on the popular Long Short-Term Memory (LSTM)  (Hochreiter & Schmidhuber,, 1997), the State-Space models (SSMs)  (Gu et al.,, 2022b; Gu & Dao,, 2023), and the Linear Recurrent Unit (LRU)  (Orvieto et al.,, 2023). All models consider conditioning on control parameters using the Feature-wise Linear Modulation (FiLM) method  (Perez et al.,, 2018) and employ the Gated Linear Unit (GLU)  (Dauphin et al.,, 2017). The architectures have a limited computational complexity, measured in the number of floating points operations per output audio sample, ensuring theoretical real-time computational capabilities on consumer-grade devices. The designed architectures also present limited intrinsic input-output audio samples latency, which we have limited to a maximum of $64$ samples ($1.33$ ms). Specifically, we refer to the intrinsic latency given by the architecture constraints, neglecting the latency given by the external DSP processing. This latency is due to the lag between the time index of the input and output samples.
@@ 2 Background
Data-driven modeling is used when the inner circuitry of the audio effect is unknown. It treats the model as a mapping between input-output data: input-output measurements are used to approximate the system response. The procedure involves three main components: the dataset, candidate modeling techniques, and the validation process. The quality of the resulting model heavily depends on the selection of input signals, configuration of control parameters, and data collection. Data-driven approaches rely on measurements that are static representations of the system behavior, leading to greater difficulty in accurately emulating its dynamic aspects.
Common examples of data-driven approaches include the block-oriented Hammerstein model, the Wiener model, and their combinations (Schoukens & Tiels,, 2017). In the Wiener model, a linear dynamic system is followed by a static nonlinearity, whereas in the Hammerstein model, the static nonlinearity precedes the linear dynamic system. These models offer a straightforward representation and are derived by estimating the model coefficients from real data using error-minimization techniques. A proposed method to classify whether a Wiener or a Hammerstein model should rather model the nonlinear system is in (Primavera et al.,, 2018). Recent examples are for guitar distortion pedal  (Eichas et al.,, 2015)  (Köper & Holters,, 2020)  (Darabundit et al.,, 2019), amplifier  (Jr & Bodmann,, 2016)  (Eichas et al.,, 2017), distortion circuit  (Eichas & Zölzer,, 2016).
A more recent data-driven technique involves utilizing artificial neural networks (ANNs). ANNs consist of interconnected units called neurons, where each connection can transmit a signal. Each neuron processes these signals and can send them to other connected neurons. The transmitted signal is represented by a real number, and the neuron’s output is determined by applying a nonlinear function to the sum of its weighted inputs. Training ANNs involve optimizing the network’s parameters to minimize the error between predicted and actual target values in a dataset. Gradient-based techniques, like backpropagation, are commonly used to optimize the network’s weights.
@@ 2.1 Neural Networks for Virtual Analog Modeling
In the last two decades, ANNs have been extensively used for modeling analog audio devices. In particular, traditional ML techniques, initially developed for different purposes, have been investigated to create models that emulate audio effects when taking and processing raw audio samples, such as traditional DSP techniques.
@@
In an early attempt using machine learning for audio effect modeling  (Mendoza,, 2005), a multilayer feedforward network was utilized to learn a high-pass filter and the Ibanez Tube-Screamer distortion effect. The high-pass filter training led to unsatisfactory results, while the distortion effect demonstrated the ability of the network to learn distortion effects.
Other attempts used convolutional-based networks to model the preamplifier circuit from the Fender Bassman 56F-A vacuum-tube amplifier  (Damskägg et al.,, 2019a). The network consists of a stack of dilated causal convolution and fully connected layers. The dataset was simulated with different values of the gain, which is a variable resistor, and used to condition the model. The conditioning is realized in the convolutional layers and using the gated activation  (Oord et al.,, 2016). The work shows that CNNs accurately emulate the reference device better than a fully connected feedforward network and can respond to user control changes.
Using the same architecture, the Ibanez Tube Screamer, the Boss DS-1, and the Electro-Harmonix Big Muff Pi distortion pedals have been modeled as well  (Damskägg et al.,, 2019b). Here, the authors investigated more configurations, considering different dilation patterns and activation functions. In this work, the conditioning is not considered.
@@
Similarly to a fully connected feedforward network, in which networks’ ability to track the past information depends on the amount of input received, the convolutional network performance depends on the receptive field; a longer receptive field requires a greater number of layers. To overcome this problem, Temporal Convolutional networks (TCN) have been proposed for optical dynamic range compressors such as Teletronix LA-2A Leveling Amplifier (Steinmetz et al.,, 2022), which present longer time-variant characteristics than distortion units. TCNs’ strength consists of dilated convolutional networks with rapidly growing dilation factors ($2^{2}$ and $02^{1}$ 0). Increasing the dilation factors increases, in turn, the receptive field and the ability of the networks when dealing with long dependencies while, in turn, not increasing the number of layers. The authors compared TCN networks with presenting receptive fields of $100$, $300$, and $1000$ ms, respectively. The model with a receptive field of $300$ ms provided the best performances. TCNs provide accurate and efficient implementation with respect to computational complexity. On the other hand, convolutional architectures present an intrinsic latency equal to the length of the receptive field, which is detrimental for live audio applications.
@@
A model that combines different deep learning architectures was investigated in  (Ramírez et al.,, 2019). In this case, the model is divided into three parts: an adaptive front-end, latent space, and synthesis back-end. The front-end performs time-domain convolutions with the raw audio to map it into a latent space. The adaptive front-end consists of a convolutional encoder. The latent-space DNN modifies the latent representation of the input audio into a new latent representation, which is fed into the synthesis back-end. The synthesis back-end consists of an unpooling layer, a deep neural network with smooth adaptive activation functions (DNN-SAAF), and a single CNN layer. The SAAF  (Hou et al.,, 2017) is used as the nonlinearity for the last layer. Distortion, overdrive, and EQ effects have been modeled using this method.
In  (Ramírez et al.,, 2020a), the previously described architecture is further explored and compared with different variations and with the CNN architecture based on  (Damskägg et al.,, 2019a). The variations mainly consist of incorporating Bi-LSTMs or CNNs into the latent space. In this work, the target devices were Universal Audio Vacuum-Tube Preamplifier 610-B, Universal Audio Transistor-Based Limiter Amplifier 1176LN, and 145 Leslie Speaker Cabinet. The highest score in the listening test was obtained by the variation incorporating Bi-LSTM. Another variation based on Bi-LSTM and SAAF as latent-space modules is used for plate and spring reverberators  (Ramírez et al.,, 2020b). These architectures represent a general-purpose deep learning framework for modeling audio effects, which has been shown to be capable of accurately modeling different types of effects. On the other hand, the models are large networks with high computational complexity, leading to difficulty when considering efficient implementation. In addition, using the Bi-LSTM layer makes the architecture non-causal, not allowing real-time applications. This architecture also lacks conditioning possibility and focuses on modeling a static representation of the devices (i.e., fixed parameters).
The last group of machine-learning black-box modeling methods we review in this section are those based on RNNs. These have been applied to model a vacuum-tube amplifier  (Covert & Livingston,, 2013), where authors investigated different networks and input sizes in two scenarios: slightly distorted and overdriven signals. The networks’ input included a variable number of audio samples. The network is autoregressive: it generates the output sample at the current time step using current and past input samples, but also output samples generated at previous time steps. The work presents preliminary experiments, and although the models resulted in poor performance, the work demonstrates the feasibility of modeling vacuum-tube amplifiers with RNNs.
RNNs, in particular LSTMs, have also been compared to CNNs when applied to the Ibanez Tube Screamer, the Boss DS-1, and the Electro-Harmonix Big Muff Pi distortion pedals  (Wright et al.,, 2019). The architecture consists of one RNN layer and one linear fully connected layer as an output layer with one unit. LSTM and GRU are compared as RNN layers varying the number of units. The input signal is extended to include an additional value representing the user control’s parameter. The results show that RNNs achieve similar accuracy to the CNN model while requiring significantly less computational cost and latency since requiring only the current input sample to predict the current output. LSTM performed better than GRU networks, and the error decreased when the number of units was increased.
@@
Another RNN-based model utilized for the Klon Centaur guitar pedal circuit is presented in  (Chowdhury,, 2020), which is also compared with a WDF model of the same circuit. ML implementation resulted in better computational performance than the WDF model but sounded very similar, although the ML implementation slightly damped high frequencies. The architecture consists of a GRU layer and one linear fully connected layer with one unit as the output layer. Also, here, a single variable control parameter of the audio effect is added as an extra input to condition the inference of the networks.
A delay-based effect is investigated in  (Mikkonen et al.,, 2023), where a tape-based delay is modeled using GRU networks. Here, the delay trajectory is analyzed and extracted using impulse train signals. This trajectory is used to demodulate the signal before training the models or used to directly guide a delay-line built in a differentiable way and based on  (Engel et al.,, 2020) work.
RNNs rely on their internal memory mechanism to encode the history of the signal, but they can still suffer from remembering significant long-term dependencies. Indeed, LSTMs alone have not been successful for LFO delay-based effects, such as flanger and phaser. In  (Wright & Välimäki,, 2021), authors found it beneficial to use as extra input the LFO signal representing the frequency response of a time-varying system over time. The LFO signal is first extracted by analyzing the recording and then used to help the LSTM during the learning process. The extra input has shown to be beneficial to learning the LFO-based effect and can be freely controlled after model training, although in the study is considered only static parameter values and not LFO signal changing depth and width over time.
A sequence-to-sequence-based model, based on LSTMs, was designed and proposed for the case of another optical compressor, such as the TubeTech CL 1B compressor  (Simionato & Fasciani,, 2022), to help LSTMs’ memory. The network is based on two LSTM layers: one acts as an encoder, and the other acts as a decoder. The first takes a segment of past input samples together with the controls’ parameters, and the decoder takes the current input sample (in case of single output) or samples (in case of multi-sample output). The encoder, when processing the input, computes the internal states. These states are used as initial states for the second LSTM. These states represent a summary of the encoder input, which helps the decoder to infer the output. The sequence-to-sequence-based model is compared to LSTMs and fully connected networks, resulting in significantly better accuracy. The authors initially considered two control parameters, threshold, and ratio, while attack and release parameters were fixed. In  (Simionato & Fasciani,, 2023), the latter parameters were included. The model’s encoder was replaced with a convolutional layer, and the model was compared to TCNs. Sequence-to-sequence-based models allow for more responsiveness and interactivity but at the expense of an increasing computational cost.
Other works consider hybrid approaches. In  (Parker et al.,, 2019), fully connected neural networks exploit measurements within the circuits. Here, the network is embedded within a discrete-time state-space system and explored for diode-based guitar distortion circuits and the lowpass filter of the Korg MS-20 synthesizer. The method required measurements from the devices’ circuits, which are not always accessible. Continuing towards gray box approaches,  (Kuznetsov et al.,, 2020) is based on the Differentiable Digital Signal Processing (DDSP) framework and applies this concept to RC lowpass filter. Subsequently, the idea is applied to the parametric tone section of a guitar amplifier and a nonlinear guitar overdrive circuit in (Esqueda et al.,, 2021). The technique learns the parameters of a circuit model from measurements of real devices using a Wiener-Hammerstein model consisting of two IIR filters and a feedforward artificial neural network. The IIR filters are, in this way, trainable and derived from the analysis of the circuit. This approach is also used in (Nercessian et al.,, 2021). In this case, a cascade of biquad filters is trained, employing a Fourier transform, to model the BOSS MT-2 pedal distortion pedal. The model includes four tunable parameters. Differentiable Digital Filters have been applied for the case of a diode clipper circuit (Chowdhury & Clarke,, 2022), showing higher accuracy and similar computation time when compared to traditional physics-based models. In (Wilczek et al.,, 2022), the authors proposed to learn ordinary differential equations governing the first-order and the second-order diode clipper using fully connected layers and RNNs. The formulation similarly uses the discrete-time state-space as in (Parker et al.,, 2019). Neural networks have also been investigated in the case of wave digital filter formulations (Darabundit et al.,, 2022). A tube amplifier circuit is simulated using multiport nonlinearities inside a wave digital filter and fully connected layers to learn features from a Kirchhoff-domain dataset. Similarly, in  (Miklanek et al.,, 2023), authors combine recurrent neural networks (RNNs) with the Kirchhoff nodal analysis to create a grey-box guitar amplifier model. Lastly, an example of gray box modeling in the case of the LA-2A leveling amplifier is (Wright & Välimäki,, 2022), where it is modeled starting from the structure of a traditional digital compressor but using fully connected and RNN to predict the compressor’s parameters.

### v6

@@ Abstract
Artificial neural networks are a promising technique for virtual analog modeling, having shown particular success in emulating distortion circuits. Despite their potential, enhancements are needed to enable effect parameters to influence the network’s response and to achieve a low-latency output. While hybrid solutions, which incorporate both analytical and black-box techniques, offer certain advantages, black-box approaches, such as neural networks, can be preferable in contexts where rapid deployment, simplicity, or adaptability are required, and where understanding the internal mechanisms of the system is less critical. In this article, we explore the application of recent machine learning advancements for virtual analog modeling. In particular, we compare State-Space models and Linear Recurrent Units against the more common Long Short-Term Memory networks, with a variety of audio effects. We evaluate the performance and limitations of these models using multiple metrics, providing insights for future research and development. Our metrics aim to assess the models’ ability to accurately replicate the signal’s energy and frequency contents, with a particular focus on transients. The Feature-wise Linear Modulation method is employed to incorporate effect parameters that influence the network’s response, enabling dynamic adaptability based on specified conditions. Experimental results suggest that Long Short-Term Memory networks offer an advantage in emulating distortions and equalizers, although performance differences are sometimes subtle yet statistically significant. On the other hand, encoder-decoder configurations of Long Short-Term Memory networks and State-Space models excel in modeling saturation and compression, effectively managing the dynamic aspects inherent in these effects. However, no models effectively emulate the low-pass filter, and Linear Recurrent Units show inconsistent performance across various audio effects.
@@ 1 Introduction
The emulation of analog musical devices has become pervasive in digital products due to their distinct sound, appealing to producers and musicians. Analog circuits use components such as operational amplifiers and diodes whose non-linear behaviors contribute to creating a unique sound. Replicating these devices, called Virtual Analog (VA) modeling, is a crucial area in digital audio signal processing.
@@
There are two possible approaches for emulating analog audio effects: physics-based and data-driven. Physics-based approaches discretize existing continuous-time models. Data-driven approaches identify a discrete model directly from discrete measurement data. Machine learning (ML), as a data-driven method, has become widely used in audio modeling and is often integrated into more traditional digital signal-processing frameworks. Hybrid solutions, incorporating knowledge of the modeling tasks, also exist and can offer advantages, but data-driven approaches can still be appealing in some contexts. While the former may not always be accurate enough, the latter may not be feasible for certain tasks.
@@
Data-driven models, also called Black-box models, have been utilized for different audio modeling tasks  Ramírez et al., (2020a), but these usually involve large networks that are not suitable for low-latency and real-time applications. In addition, while machine learning modeling of nearly all types of analog effects has been investigated, they often do not provide a satisfactory response when continuously manipulating conditioning parameters, which is needed to offer users runtime control over the effect’s parameters.
Neural networks incorporating internal states and utilizing recurrent formulations have emerged as the most accurate option for modeling audio effects among various types of artificial neural networks, with Recurrent Neural Networks (RNNs) being the most recognized. A similar formulation exists in State-Space Models (SSMs). The internal states in these architectures serve as memory, rendering them particularly effective in handling scenarios with transients and time dependencies, especially when processing one or a few audio samples at a time to achieve low latency. Consequently, in this study, we include Convolutional Neural Networks (CNNs) LeCun et al., (1998)—which are inherently stateless—only when integrated with RNNs Simionato & Fasciani, (2022).
This article presents a comprehensive comparative study that systematically evaluates state-based neural network architectures in modeling various types of audio effects, each posing unique modeling challenges, and in the context of achieving low-latency, computationally sustainable, and parametric solutions. Experimental results highlight the strengths and limitations of different state-based networks in their ability to learn diverse audio effects. Specifically, we explore the learning capabilities of these architectures, investigating the similarities and differences between them. By analyzing RNNs and SSMs, which use distinct recurrence methods to compute internal states, we provide insights into their variant modeling capabilities across different contexts. While architecture-specific optimizations, such as hyperparameter tuning or layer size adjustments, were not pursued, our study prioritized limiting computational complexity and minimizing input-output latency. We focused on developing models that are suitable for implementation on consumer-grade devices for live audio applications.
@@
We designed five architectures based on the popular Long Short-Term Memory (LSTM)  Hochreiter & Schmidhuber, (1997), the SSMs  Gu et al., (2022b); Gu & Dao, (2023), and the Linear Recurrent Unit (LRU)  Orvieto et al., (2023). All models consider conditioning on control parameters using the Feature-wise Linear Modulation (FiLM) method  Perez et al., (2018) and employ the Gated Linear Unit (GLU)  Dauphin et al., (2017). The architectures are engineered to operate with a computational load of less than $100$ MFLOPS (Million Floating Point Operations Per Second) when processing audio at $48$ kHz. This constraint facilitates the simultaneous operation of multiple software audio effects on consumer-grade personal computers and supports independent execution on bare-metal embedded systems. Additionally, the architectures are designed to exhibit minimal inherent input-output latency for audio samples. Specifically, we refer to the intrinsic latency given by the architecture constraints, neglecting the latency given by the external Digital Signal Processor (DSP) processing. This latency is due to the lag between the time index of the input and output samples.
@@ 2 Background
Data-driven modeling is typically used when the internal circuitry of an audio effect is unknown, treating the model as a mapping of input-output data to approximate the system’s response. This process involves three main components: the dataset, modeling techniques, and validation methods. The model’s quality is heavily dependent on the selection of input signals, control parameter configurations, and data collection strategies. However, because data-driven approaches often rely on static measurements, accurately capturing dynamic behavior presents a notable challenge.
Common examples of data-driven approaches include the block-oriented Hammerstein model, the Wiener model, and their combinations  Schoukens & Tiels, (2017). Recent examples of their application focus on guitar distortion pedal  Eichas et al., (2015)  Köper & Holters, (2020)  Darabundit et al., (2019), amplifier  Jr & Bodmann, (2016)  Eichas et al., (2017), distortion circuit  Eichas & Zölzer, (2016).
More recent data-driven techniques leverage artificial neural networks (ANNs), comprising interconnected nonlinear units known as neurons. Each neuron propagates signals through its connections, processes these signals, and has the ability to transmit the processed signals to other connected neurons. Training ANNs involves optimizing the network’s parameters to minimize the error between predicted and actual target values in a dataset. Gradient-based techniques, like backpropagation, are commonly used to optimize the network’s weights.
@@ 2.1 Neural Networks for Virtual Analog Modeling
In the last two decades, ANNs have been extensively used for modeling analog audio devices. In particular, traditional ML techniques have been investigated to create models that emulate audio effects processing raw audio samples, such as traditional DSP techniques.
@@
In an early attempt using ML for audio effect modeling  Mendoza, (2005), a multilayer feedforward network was utilized to learn a high-pass filter and the Ibanez Tube-Screamer distortion effect. The high-pass filter experiments led to unsatisfactory results as the model did not learn any aspect of the filtering process, while with the distortion effect, the network demonstrated the ability to replicate the signal alteration typical of such audio effects accurately.
Other studies utilized CNNs to model the preamplifier circuit of the Fender Bassman 56F-A vacuum-tube amplifier  Damskägg et al., (2019a). The proposed network features stacked dilated causal convolution and fully connected layers, trained on a synthetic dataset with varying gain values for conditioning the network through a gated activation technique  Oord et al., (2016). Results indicate that CNNs outperform fully connected networks in emulating the device and they can also adapt to changes in the effect’s control parameters. The same architecture was later applied to model the Ibanez Tube Screamer, Boss DS-1, and Electro-Harmonix Big Muff Pi distortion pedals  Damskägg et al., (2019b), exploring various configurations without considering effects parameters for conditioning the network.
@@
Similarly to fully connected networks, CNNs’ ability to track the input’s past information depends on their receptive field, which requires more layers when the fields are longer. Temporal Convolutional Networks (TCNs) use dilated convolutions with increasing dilation factors ($2^{2}$ and $2^{10}$), enhancing the receptive field and long dependency handling without increasing layer count. TCNs have been applied to optical dynamic range compressors, such as the Teletronix LA-2A Leveling Amplifier  Steinmetz et al., (2022), where the utilization of a $300$ ms receptive field yielded the best performance. TCNs provide accurate and computationally efficient implementations, but network architectures with convolutional layers present intrinsic input-output latency equal to the receptive field length, which can hinder live audio applications. TCNs have also been used to design Open-Amp  Wright et al., (2025), a model emulating many guitar effects. The model is conditioned on a learnable look-up table of embeddings and learns one embedding vector per effect. In this case, the length of the receptive field is not reported.
@@
A model combining various deep learning architectures was explored in  Ramírez et al., (2019). The model includes an adaptive front-end, a latent space, and a synthesis back-end. The front-end uses a convolutional encoder to perform time-domain convolutions on raw audio, mapping it into a latent space. The latent-space transforms the input audio representation, which is then processed by the synthesis back-end, consisting of an unpooling layer, a deep neural network with smooth adaptive activation functions (SAAF)  Hou et al., (2017). This approach has been applied to model distortion, overdrive, and equalizer effects.
In  Ramírez et al., (2020a), the previously described architecture was further investigated and compared with variations and the CNN-based architecture from  Damskägg et al., (2019a). Variations included incorporating Bi-LSTMs or CNNs into the latent space. The Bi-LSTM variation achieved the highest score in listening tests. Another Bi-LSTM and SAAF variation was used for plate and spring reverberators  Ramírez et al., (2020b). While these architectures effectively model various audio effects, they are large networks with high computational complexity, making efficient implementation challenging. The use of Bi-LSTMs also introduces non-causality, hindering real-time applications and interactivity with the effect’s parameters. In  Comunità et al., (2023), LSTMs are integrated to enhance intermediate data representation in TCN models, tested on fuzz and compressor effects, and the proposed architecture outperforms the baselines of LSTM and CNN. As in  Steinmetz et al., (2022), the models present long receptive fields, in this case, between $45$ ms and $2500$ ms.
Finally, RNNs have been used to model vacuum-tube amplifiers  Covert & Livingston, (2013). Although preliminary experiments showed limited accuracy, they established RNN feasibility for this modeling task. RNNs, particularly LSTMs, were also compared to CNNs in modeling the Ibanez Tube Screamer, Boss DS-1, and Electro-Harmonix Big Muff Pi distortion pedals  Wright et al., (2019). The architecture included one RNN layer and a linear output layer, with LSTM and Gated Recurrent Unit (GRU) compared by varying the number of units. The input of the network includes both the audio signal and the effect parameters. The results indicated that RNNs achieved similar accuracy to CNNs, but with lower computational costs and latency, relying only on the current input sample for output prediction. LSTM outperformed GRU, with error reduction observed as unit count increased.
@@
In  Chowdhury, (2020), another RNN-based model for the Klon Centaur guitar pedal circuit is compared with a Wave Digital Filter (WDF) model. The machine learning approach presented a lower computational cost, though the emulated sound was very similar to the WDF model, with a slight damping of high frequencies. The architecture includes a GRU layer and a linear output layer with one unit, and a single control parameter is also fed as input to the network to condition the inference.
In  Mikkonen et al., (2023), a tape-based delay effect is modeled using GRU networks, with the delay trajectory analyzed through impulse train signals. This trajectory demodulates the signal before model training or guides a differentiable delay line based on  Engel et al., (2020). In  Wright & Välimäki, (2021), the modeling of delay-based effects, such as flanger and phaser, is investigated using an extracted low-frequency oscillation (LFO) signal to aid LSTMs during training. This is provided as additional input to the network, and it is shown to improve the learning of these effects. A the same time, it allows for post-training control of the model; however, the experiments were focused on fixed effect parameters, without changes in the LFO depth and width over time.
An encoder-decoder (ED) network was utilized to model the TubeTech CL 1B compressor  Simionato & Fasciani, (2022), featuring an encoder LSTM that processes past input samples and control parameters, and a decoder LSTM handling current and more recent past input samples. The encoder’s internal states summarize the input, aiding the decoder’s output inference. This model outperformed standard LSTMs and fully connected networks in accuracy, initially modeling only threshold and ratio control parameters, while the attack and release settings were fixed. In  Simionato & Fasciani, (2023), all four compressor control parameters were modeled, and the encoder was replaced with a convolutional layer for comparison with TCNs. While ED models offer increased responsiveness and interactivity, they also incur higher computational costs.
Other studies explored hybrid modeling approaches. In  Parker et al., (2019), fully connected neural networks utilize circuit measurements, embedding the network within a discrete-time state-space system for diode-based guitar distortion circuits and the Korg MS-20 lowpass filter. This method relies on circuit measurements, which may not always be accessible. Similarly,  Peussa et al., (2021) proposes a circuit state-matching mechanism for GRU networks, training a variant of the GRU layer to match circuit states using a teaching-forcing method.
The Differentiable Digital Signal Processing (DDSP) framework for modeling an RC lowpass filter is employed in  Kuznetsov et al., (2020), later also used for the parametric tone section of a guitar amplifier and a nonlinear overdrive circuit in  Esqueda et al., (2021). This method learns circuit parameters from real device measurements using a Wiener-Hammerstein model with two trainable Infinite Impulse Response (IIR) filters and a feedforward neural network. Similarly,  Nercessian et al., (2021) trains a cascade of biquad filters via the Fourier transform for the BOSS MT-2 distortion pedal, incorporating four tunable parameters. The DDSP framework was also successfully applied to a diode clipper circuit in  Chowdhury & Clarke, (2022), demonstrating higher accuracy with comparable computation times to traditional models.
In  Wilczek et al., (2022), the authors learn ordinary differential equations governing diode clippers using fully connected layers and RNNs, aligned with the discrete-time state-space approach from  Parker et al., (2019). Investigations into WDF formulations are found in  Darabundit et al., (2022), simulating tube amplifier circuits with multiport nonlinearities and fully connected layers.  Miklanek et al., (2023) combines RNNs with Kirchhoff nodal analysis to model guitar amplifiers. A hybrid approach for the LA-2A leveling amplifier is presented in  Wright & Välimäki, (2022), modeled using a traditional digital compressor structure alongside fully connected RNNs for parameter prediction. Lastly,  Huhtala et al., (2024) proposed a Koopman-Linearised Audio Neural Network structure, which comprises two gated linear unit multilayer perceptrons and differentiable biquad filters.